please,solve this problem for me 2+4-6-3x2+(3x7)=?

Found 2 solutions by RAY100, shymeerab:
Answer by RAY100(1637)

Remember order of operations, PEMDAS, Parenthesis, Exponents, Multiply / Divide, lt to rt, and ADD / Subtract, lt to rt
.
2+4 -6 -3*2 +3*7
.
2+4 -6 -6 +21
.
6 -6 -6 +21
.
0-6 +21
.
-6 +21
.
15

Answer by shymeerab(21)

=2+4-6-(3*2)+(3*7)
=0-(6)+(21)
=-6+21
=15